Hi
A is 5km from a lighthouse on a bearing N37W. B is 12km from the same lighthouse on a bearing of S53W.
find the distance and bearing of B from A.
thanks

Found 2 solutions by Alan3354, addingup:
Answer by Alan3354(69443)

A is 5km from a lighthouse on a bearing N37W. B is 12km from the same lighthouse on a bearing of S53W.
find the distance and bearing of B from A.
--------------------
2 sides & the included angle --> Cosine Law

Answer by addingup(3677)

sqrt(5^2 + 12^2) = sqrt(25 + 144) = sqrt(169) = 13km
the direction of this 13km hypotenuse from A to B is S78W (I think that's what you'd call it, for me it's 78°SSW)